# From Optimization to Control: Quasi Policy Iteration

### v3

@@ 6. Limitations and Future Research
The main drawback of the proposed algorithms, similar to other accelerated VI schemes in the literature, is the need for safeguarding to ensure convergence. Our experiments in Section 5 showed examples of MDPs in which the safeguard is activated. Our modified implementation of Section 3.2.2 guarantees the linear convergence while removing the need for safeguarding and using backtracking instead. Another possible solution is the use of the operator splitting method introduced in [44] for policy evaluation. In this regard, we note that the proposed QPI algorithm is essentially the PI algorithm in which the policy evaluation step uses the approximation $\widetilde{P}_{k}$ in (12) instead of the true transition matrix $P_{k}$ and the cost $\widetilde{c}_{k}=c_{k}+\gamma(P_{k}-\widetilde{P}_{k})v_{k}$ instead of the true cost $c_{k}$. However, the convergence requires $\widetilde{P}_{k}$ to be close to $P_{k}$. To be precise, a sufficient condition is $\|P_{k}-\widetilde{P}_{k}\|_{\infty}\leq 1-\gamma$ [44, Thm. 1], which is difficult to achieve for a low-rank approximation $\widetilde{P}_{k}$ of $P_{k}$.
Another limitation of the current work is the lack of a theoretical guarantee for the empirically observed improvement in the convergence rate, particularly for the model-based QPI algorithm. The main difficulty to be addressed is the fact that the linear constraints in (12) are not the standard secant conditions used in QNMs. The modified implementation of Section 3.2.3 partially addresses this issue by including the secant conditions in its approximation. An interesting result, however, is the establishment of a local superlinear convergence rate for QPI without imposing the secant condition. Another possibility is to use the results for Anderson acceleration in [16] to establish an improved linear rate for convergence. To that end, similar to what is done in [48], one needs to use a smoothed version of the Bellman operator, e.g., by replacing the max operation with a soft-max operation in the Bellman operator.
@@ Appendix A MDPs of the numerical simulations
Garnet MDP. The considered Garnet MDP [3] is generated randomly with $n=50$ states, $m=5$ actions, and the branching parameter $n_{b}=10$. For each state-action pair $(s,a)$, we first form the set of reachable next states $\{s_{1}^{+},\ldots,s_{n_{b}}^{+}\}$ chosen uniformly at random from the state space $\{1,\ldots,n\}$. Then, the corresponding probabilities are formed by choosing the points $p_{i}\in[0,1],i=1,\ldots,n_{b}-1$, uniformly at random, and setting $\mathds{P}(s_{i}^{+}|s,a)=p_{i}-p_{i-1}$ with $p_{0}=0$ and $p_{n_{b}}=1$. The stage cost $c(s,a)$ for each state-action pair $(s,a)$ is also chosen uniformly at random from the interval $[0,1]$.
Healthcare MDP. The considered Healthcare MDP is borrowed from [22]. The MDP has 6 states corresponding to the deteriorating health condition of a patient with the last state $n=6$ being an absorbing state representing the mortality terminal state. For each of the first five states, one can choose three inputs $m\in\{1,2,3\}$ corresponding to increasing levels of drug dosage for treatment. The goal is to minimize the invasiveness of the treatment while avoiding the terminal state. For the transition probabilities, we refer the reader to [22, Fig. D.1]. The cost function is chosen to be $c(n,m)=\sum_{n^{+}=1}^{6}n^{+}\mathds{P}(n^{+}|n,m)+m$ for each $n\in\{1,2,3,4,5\}$ and $m\in\{1,2,3\}$ and $c(6,1)=50$.
Graph MDP. The considered Graph MDP is borrowed from [15]. The MDP has 18 state-action pairs in total and corresponds to a simple path-finding problem. We refer the reader to [15, Sec. 3] for the description of the MDP.
@@ Appendix B Accelerated VI and QL algorithms
(1) Nesterov accelerated VI [22] – with $v_{-1}=v_{0}$:
@@
(2) Anderson accelerated VI [19] – with $v_{-1}=v_{0}$:
@@
(3) Speedy QL [20] – with $q_{-1}=q_{0}$, $\alpha_{k}=\frac{1}{k+1}$:
@@
(4) Zap QL [15] – with $D_{-1}=\boldsymbol{0}$, $\alpha_{k}=\beta_{k}=\frac{1}{k+1}$:

### v4

@@ 6. Limitations and Future Research
The main drawback of the proposed algorithms, similar to other accelerated VI schemes in the literature, is the need for safeguarding to ensure convergence. Our experiments in Section 5 showed examples of MDPs in which the safeguard is activated. Our modified implementation of Section 3.2.2 guarantees the linear convergence while removing the need for safeguarding and using backtracking instead. Another possible solution is the use of the operator splitting method introduced in [NEURIPS2022_fa809df3] for policy evaluation. In this regard, we note that the proposed QPI algorithm is essentially the PI algorithm in which the policy evaluation step uses the approximation $\widetilde{P}_{k}$ in (12) instead of the true transition matrix $P_{k}$ and the cost $\widetilde{c}_{k}=c_{k}+\gamma(P_{k}-\widetilde{P}_{k})v_{k}$ instead of the true cost $c_{k}$. However, the convergence requires $\widetilde{P}_{k}$ to be close to $P_{k}$. To be precise, a sufficient condition is $\|P_{k}-\widetilde{P}_{k}\|_{\infty}\leq 1-\gamma$ [NEURIPS2022_fa809df3, Thm. 1], which is difficult to achieve for a low-rank approximation $\widetilde{P}_{k}$ of $P_{k}$.
Another limitation of the current work is the lack of a theoretical guarantee for the empirically observed improvement in the convergence rate, particularly for the model-based QPI algorithm. The main difficulty to be addressed is the fact that the linear constraints in (12) are not the standard secant conditions used in QNMs. The modified implementation of Section 3.2.3 partially addresses this issue by including the secant conditions in its approximation. An interesting result, however, is the establishment of a local superlinear convergence rate for QPI without imposing the secant condition. Another possibility is to use the results for Anderson acceleration in [Evans2020] to establish an improved linear rate for convergence. To that end, similar to what is done in [sun2021damped], one needs to use a smoothed version of the Bellman operator, e.g., by replacing the max operation with a soft-max operation in the Bellman operator.
@@ Appendix A MDPs of the numerical simulations
Garnet MDP. The considered Garnet MDP [archibald1995generation] is generated randomly with $n=50$ states, $m=5$ actions, and the branching parameter $n_{b}=10$. For each state-action pair $(s,a)$, we first form the set of reachable next states $\{s_{1}^{+},\ldots,s_{n_{b}}^{+}\}$ chosen uniformly at random from the state space $\{1,\ldots,n\}$. Then, the corresponding probabilities are formed by choosing the points $p_{i}\in[0,1],i=1,\ldots,n_{b}-1$, uniformly at random, and setting $\mathds{P}(s_{i}^{+}|s,a)=p_{i}-p_{i-1}$ with $p_{0}=0$ and $p_{n_{b}}=1$. The stage cost $c(s,a)$ for each state-action pair $(s,a)$ is also chosen uniformly at random from the interval $[0,1]$.
Healthcare MDP. The considered Healthcare MDP is borrowed from [goyal2019first]. The MDP has 6 states corresponding to the deteriorating health condition of a patient with the last state $n=6$ being an absorbing state representing the mortality terminal state. For each of the first five states, one can choose three inputs $m\in\{1,2,3\}$ corresponding to increasing levels of drug dosage for treatment. The goal is to minimize the invasiveness of the treatment while avoiding the terminal state. For the transition probabilities, we refer the reader to [goyal2019first, Fig. D.1]. The cost function is chosen to be $c(n,m)=\sum_{n^{+}=1}^{6}n^{+}\mathds{P}(n^{+}|n,m)+m$ for each $n\in\{1,2,3,4,5\}$ and $m\in\{1,2,3\}$ and $c(6,1)=50$.
Graph MDP. The considered Graph MDP is borrowed from [devraj2017zap]. The MDP has 18 state-action pairs in total and corresponds to a simple path-finding problem. We refer the reader to [devraj2017zap, Sec. 3] for the description of the MDP.
@@ Appendix B Accelerated VI and QL algorithms
(1) Nesterov accelerated VI [goyal2019first] – with $v_{-1}=v_{0}$:
@@
(2) Anderson accelerated VI [geist2018anderson] – with $v_{-1}=v_{0}$:
@@
(3) Speedy QL [ghavamzadeh2011speedy] – with $q_{-1}=q_{0}$, $\alpha_{k}=\frac{1}{k+1}$:
@@
(4) Zap QL [devraj2017zap] – with $D_{-1}=\boldsymbol{0}$, $\alpha_{k}=\beta_{k}=\frac{1}{k+1}$: